# From EHRs to Patient Pathways: Scalable Modeling of Longitudinal Health Trajectories with LLMs

### v1

@@ Abstract
Healthcare systems face significant challenges in managing and interpreting vast, heterogeneous patient data for personalized care. Existing approaches often focus on narrow use cases with a limited feature space, overlooking the complex, longitudinal interactions needed for a holistic understanding of patient health. In this work, we propose a novel approach to patient pathway modeling by transforming diverse electronic health record (EHR) data into a structured representation and designing a holistic pathway prediction model, EHR2Path, optimized to predict future health trajectories. Further, we introduce a novel summary mechanism that embeds long-term temporal context into topic-specific summary tokens, improving performance over text-only models, while being much more token-efficient. EHR2Path demonstrates strong performance in both next time-step prediction and longitudinal simulation, outperforming competitive baselines. It enables detailed simulations of patient trajectories, inherently targeting diverse evaluation tasks, such as forecasting vital signs, lab test results, or length-of-stay, opening a path towards predictive and personalized healthcare.
@@ 1 Introduction
Healthcare systems face increasing challenges in managing vast amounts of patient data while striving for personalized, efficient care rajkomar2018scalable ; jensen2012mining . The data collected and stored in healthcare has reached a scale where human interpretation alone is insufficient, necessitating the support of intelligent computational systems topol2019high . So far, while the progress in computer-aided diagnosis has enhanced diagnostic accuracy and efficiency khalifa2024ai , most predictive models compartmentalize patient information, failing to capture the complex interactions of health factors over time rajkomar2018scalable ; miotto2018deep . Recent efforts in Electronic Health Record (EHR) modeling have integrated structured and unstructured data, including demographics, textual reports, and time-series events yang2023machine ; jensen2012mining ; mcdermott2021EHRbenchmark ; pellegrini2023unsupervised ; lovon2024revisiting ; kraljevic2024foresight . However, these studies often use a narrow set of features, often rely on complex data curation, and evaluate on limited tasks, leaving the potential for holistically modeling patient pathways largely unexplored.
@@
Patient pathways richter2019understanding capture health trajectories at scales ranging from long-term, spanning a patient’s lifetime, tackling e.g. chronic disease progression, to short-term, focused on specific healthcare episodes like hospital stays. Owing to data availability, this work focuses on short-term hospital trajectories, as visualized in Figure 1, though our approach naturally extends to long-term pathways. A comprehensive patient pathway model that can simulate future patient trajectories, could anticipate deteriorating vital signs, forecast adverse events in order to preemptively react to them, or compare likely outcomes under different treatment plans in silico, supporting proactive, personalized care.
Recent advancements in multimodal large language models (MLLMs), have demonstrated strong capabilities in processing complex, multi-modal data anil2023palm ; touvron2023llama ; openai2023gpt4 . In medicine, LLMs have excelled in tasks such as medical exam solving singhal2023towards ; toma2023clinical , literature comprehension singhal2023large , and conversational diagnosis li2023chatdoctor ; toma2023clinical ; pellegrini2023radialog . Unlike AI models tailored for specific data types, LLMs can unify diverse inputs by transforming them into compatible representations, whether as textual, numerical, or categorical sequences, preserving contextual relationships.
We believe that accurately modeling patient pathways demands a paradigm shift to models capable of integrating all available patient data over extended time periods. Towards this, we introduce EHR2Path, the first holistic pathway foundation model trained on large-scale EHR data, covering a patient’s full stay in the hospital including emergency department, hospital, and intensive care unit (ICU) data. We transform a wide range of highly heterogeneous inputs into a unified, structured text representation that is effectively processed by building upon the knowledge and capabilities of LLMs. To capture truly long-term trajectories, we propose a novel summary mechanism that compresses past context into topic-specific “summary tokens,” enabling up to 625× more context. Finally, we demonstrate two complementary fine-tuning strategies to specialize EHR2Path to specific tasks.
We demonstrate the strong performance of EHR2Path in patient trajectory forecasting and outcome prediction on the MIMIC-IV dataset johnson2023mimic , showing improved results over baselines in pathway and outcome prediction and the effectiveness of our proposed summary method. Our work establishes both a scalable methodology and a baseline for holistic patient pathway prediction, shifting the focus from isolated outcome prediction to the comprehensive modeling of patient trajectories, thereby advancing personalized and predictive healthcare.
@@ 2 Related Work
The integration of heterogeneous EHR data into machine learning models has been a significant research focus, particularly with the adoption of sequence models like transformers (vaswani2017attention, ). Existing methods fall into two categories: Outcome Prediction Models, that target specific patient outcomes, such as mortality, length-of-stay or diagnosis prediction, and Patient Timeline Prediction Models, which forecast full health trajectories but remain under-explored.
@@
Outcome Prediction Methods   Most EHR models predict outcomes using structured data like medical codes. Models such as BEHRT li2020behrt , Med-BERT rasmy2021med , CEHR-BERT pang2021cehr , TransformEHR yang2023transformehr have enhanced accuracy with temporal embeddings, artificial time tokens and multi-task learning. EHRSHOT wornow2023ehrshot introduced a benchmark integrating structured demographics and various coded events. However, their reliance on structured codes limits capturing multi-modal EHR complexity. While some other works include additional numerical or categorical data li2022hi ; pellegrini2023unsupervised ; lovon2024revisiting , they remain constrained to a predefined and rather narrow set of features. Another major challenge in EHR modeling is Long-Context Integration, due to EHR’s time span and volume. Some outcome prediction models like Hi-BEHRT li2022hi , EhrMamba fallahpour2024ehrmamba , and CONTEXT wornow2024context extend token limits with hierarchical or sub-quadratic architectures, while models like REMed kim2023general and EMERGE zhu2024emerge propose retrieval-augmented approaches for selecting relevant features. Nevertheless, these models are restricted to outcome prediction.
@@
Patient Timeline Prediction Methods   In contrast to the prevalent outcome prediction methods, patient timeline prediction seeks to model and predict the entire sequence of a patient’s health events, offering a more holistic view of health trajectories. While this area has received less attention, some recent works have begun to explore this direction. For instance, Foresight (kraljevic2024foresight, ) predicts biomedical concepts from clinical notes, converting text into structured sequences using the SNOMED ontology (el2018snomed, ). However, it excludes time-series data, such as measurements, and does not assess long-term simulation capabilities. ETHOS (renc2024zero, ) models structured EHR data from MIMIC-IV, generating tokenized Patient Health Timelines (PHTs). Yet, it omits unstructured notes and ICU chart events, key sources of patient information, and is limited to a 2,048-token input window. Further, while it proposes a simulation-based approach, it only evaluates predicting outcomes like mortality and readmission. While these models advance pathway forecasting, their limitations to fully integrate diverse EHR data types, and their lack in validation on trajectory prediction, restrict their scope.
@@
Our work shifts the focus to forecasting and evaluating on entire health pathways, moving beyond isolated outcome prediction towards a comprehensive understanding of patient development. Unlike prior methods, we comprehensively integrate all available structured and unstructured EHR data and leverage LLMs to understand this heterogeneous and noisy information in its native form. To address long-context challenges, we introduce a novel summary mechanism that enables efficient modeling over extended time horizons and complex patient histories.
@@ 3 Scalable Modeling of Health Trajectories
We propose a novel approach to predict longitudinal patient health trajectories by integrating heterogeneous EHR data into a structured, language-based representation. Using an LLM backbone, we leverage its contextual understanding to model complex health data. To capture extended time horizons, we introduce a Masked Summarization Bottleneck that efficiently compresses longitudinal data. Additionally, we introduce two fine-tuning strategies for specializing EHR2Path to specific tasks. Our method enables comprehensive trajectory simulation, supporting holistic patient modeling.
@@ 3.1 Data Representation
EHR data is inherently heterogeneous, including static and temporal data, numerical and categorical features, and unstructured text from various clinical settings. We develop a unified textual representation of this data preserving its original form, avoiding heavy pre-processing to embrace real-world noise and incompleteness. This enhances robustness and supports diverse inputs without feature-specific adjustments, thereby improving scalability and real-world applicability. Our method supports diverse data types and uses natural language for feature names and values, instead of medical codes, enabling semantic interpretation by an LLM. An exemplary data sample is shown in Appendix C.
@@
[…]
@@
State Attributes and Diagnosis Codes   We incorporate state attributes to mark significant changes in the patient’s trajectory. These include events such as discharge from the ED, admission to and discharge from the ICU and Hospital, as well as death. At the final time step $T_{p}$ of a stay, ICD categories slee1978international corresponding to the patient’s diagnosis are added.
Input and Output Preparation   Model input and output are constructed by generating textual representations following the hierarchical structure defined above. The input includes all available data for a time window between $t-w$ up to the current time-point $t$, where $w$ denotes the input window size. To integrate longer time windows, we introduce a Masked Summarization Bottleneck, which condenses the entire patient history into one or multiple embedded representations for each category consisting of $N$ summary tokens each, allowing a window size of $t$. The output is formatted similarly but exclusively contains data recorded at the subsequent time-point $t+1$. To enhance efficiency, the output is sparse, including only the features that were actively recorded at $t+1$.
@@ 3.2 Patient Trajectory Prediction
Building upon the structured, language-based EHR representation outlined in Section 3.1, we now detail our approach for predicting future patient states and clinical pathways.
@@ 3.2.1 Model Architecture and Training
EHR2Path-Text (E2P-T): Processes a text representation of static patient data and events within the most recent $w$ hours (e.g., $[t-w,t]$), capturing detailed recent observations in a structured format.
EHR2Path-Summ (E2P-S): Relies on summary embeddings of the entire patient history, generated via the Masked Summarization Bottleneck (Section 3.2.2), efficiently integrating long-term context.
EHR2Path-Summ+Text (E2P-S+T): Combines the text representation of recent events and the full history summary, balancing detailed short-term data with comprehensive longitudinal context.
Appendix D provides detailed implementation and training details.
@@ 3.2.2 Masked Summarization Bottleneck
For E2P-S and E2P-S+T, we propose a novel Masked Summarization Bottleneck that produces compact, task-relevant representations for each clinical unit $u$ and data category $c$ to handle extensive patient histories. Formally, let $\mathbf{X}_{t}=\{x_{1},x_{2},\dots,x_{n}\}$ be the $n$ token long description of a section in the patient’s record at time $t$. We append $m$ summary tokens $\mathbf{S}=\{s_{1},\dots,s_{m}\}$ and $o$ output tokens $\mathbf{Y}=\{y_{1},\dots,y_{o}\}$ describing the values at $t+1$ for the current section, forming the sequence $\mathbf{X}^{\text{full}}_{t}=\{\mathbf{X}_{t},\mathbf{S},\mathbf{Y}\}$. A custom attention mask $\mathbf{M}$ constrains how tokens $i$ can attend to tokens $j$:
@@
As visualized in  Figure 3, this ensures that output tokens $\{y_{1},\dots,y_{o}\}$ can only attend to the summary tokens $\{s_{1},\dots,s_{m}\}$ and preceding output tokens, creating a bottleneck. Following the information bottleneck principle tishby99information , our approach derives a compact representation $\mathbf{S}$ that maximizes mutual information $I(\mathbf{S};\mathbf{z}_{t+1})$ with the next state $\mathbf{z}_{t+1}$ while constraining $I(\mathbf{S};\mathbf{Z}_{1:t})$ with the patient history $\mathbf{Z}_{1:t}$ subject to the capacity $m$. Consequently, the Summarization Module is driven to compress the most relevant information to predict future states into the summary tokens, while discarding uninformative or redundant details. During inference, the summary tokens provide a lightweight embedding of the entire patient history for each data category $c$ in $D_{p,t}$. By handling each section separately, this allows us to in most cases include the entire patient history. At inference, the input tokens $\{x_{1},\dots,x_{n}\}$ for each category $c$ are combined with the summary tokens $\{s_{1},\dots,s_{m}\}$, and a single forward pass computes their hidden states $\mathbf{H}=\{\mathbf{h}_{1},\dots,\mathbf{h}_{n+m}\}$. The hidden states of the summary tokens, $\mathbf{H}^{\text{summary}}=\{\mathbf{h}_{n+1},\dots,\mathbf{h}_{n+m}\}$ are the final output of the Summary Model. This enables efficient summarization of all sections in $D_{p,t}$ without separate models, thus reducing computational overhead.
Figure 3: Masked Summarization Bottleneck. Input tokens encode past observations, while summary tokens ($<$SUM$>$) compress key information. A custom attention mask ensures outputs attend only to summaries, forcing the model to encode relevant patient data into a compact representation.
Unlike traditional auto-encoder approaches that reconstruct every detail of the input, we optimize summary tokens for forecasting the patient’s next state. This task-focused approach avoids the inefficiencies associated with generic reconstructions and produces more effective representations for patient pathway prediction.

### v2

# EHR2Path: Scalable Modeling of Longitudinal Patient Pathways from Multimodal Electronic Health Records

@@ Abstract
Forecasting how a patient’s condition is likely to evolve, including possible deterioration, recovery, treatment needs, and care transitions, could support more proactive and personalized care, but requires modeling heterogeneous and longitudinal electronic health record (EHR) data. Yet, existing approaches typically focus on isolated prediction tasks, narrow feature spaces, or short context windows, limiting their ability to model full patient pathways. To address this gap, we introduce EHR2Path, a multimodal framework for forecasting and simulating full in-hospital patient pathways from routine EHRs. EHR2Path converts diverse clinical inputs into a unified temporal representation, enabling modeling of a substantially broader set of patient information, including radiology reports, physician notes, vital signs, medication and laboratory patterns, and dense bedside charting. To support long clinical histories and broad feature spaces, we introduce a Masked Summarization Bottleneck that compresses long-term history into compact, task-optimized summary tokens while preserving recent context, improving both performance and token efficiency. In retrospective experiments on MIMIC-IV, EHR2Path enables next-step pathway forecasting and iterative simulation of complete in-hospital trajectories, while outperforming strong baselines on directly comparable tasks. These results establish a foundation for scalable pathway-level modeling from routine EHRs supporting anticipatory clinical decision-making. Our code is available at https://github.com/ChantalMP/EHR2Path.
@@ 1 Introduction
Anticipating how patients evolve during a hospital stay remains a central challenge in data-driven clinical decision support. Hospital care depends on repeated decisions about monitoring, treatments, transfer, and discharge, all of which require anticipating how an individual patient’s condition is likely to evolve. EHRs capture structured diagnosis and treatment codes, continuous physiologic measurements, medication administrations, bedside observations and free-text notes, offering a detailed but highly heterogeneous view of patient trajectories. Patient pathways (Richter & Schlieter, 2019) span multiple temporal scales, from long-term disease progression over a lifetime to short-term trajectories within a single healthcare episode. Here, we focus on full in-hospital patient pathways, as visualized in fig.˜1, which are documented in detail in EHRs and represent a critical window for clinical intervention. Unlike models that predict a single downstream outcome, a comprehensive in-hospital patient pathway model could forecast how a patient’s full state is likely to evolve over time, enabling anticipation of deteriorating vital signs, adverse events, treatment needs and care transitions, and thereby supporting more proactive and personalized care. Yet translating these heterogeneous records into pathway-level forecasts remains difficult. To address this challenge, we propose EHR2Path, a method for forecasting and simulating full in-hospital patient pathways from routine EHRs.
@@
Substantial progress has been made in machine learning for EHRs, with a large body of work focusing on outcome prediction, targeting individual downstream endpoints such as mortality, length of stay, diagnosis prediction, or specific laboratory values. Another line of work moves toward trajectory modeling, aiming to represent or generate sequences of patient events over time. While both directions have produced important advances, current methods often operate on restricted subsets of the EHR, limited context windows, or narrowly defined tasks. As a result, they typically do not model the full multimodal, longitudinal patient state at pathway level, and they rarely evaluate iterative simulation of complete in-hospital trajectories.
Recent advances in large language models (LLMs) provide a promising foundation for pathway-level EHR modeling. Their key advantage in this setting is the ability to represent highly heterogeneous clinical information, including free text, numerical measurements, and categorical events, within a unified autoregressive modeling framework while preserving temporal and contextual structure. By serializing diverse clinical signals into a shared sequence space, LLMs offer a flexible alternative to modality-specific pipelines and narrowly scoped task models. Prior work has shown that such models can learn rich representations from clinical text (Yang et al., 2022) and can support unified modeling of diverse EHR signals (Hur et al., 2023; McDermott et al., 2023; Makarov et al., 2025). However, directly applying LLMs to full hospital trajectories remains challenging, since clinically relevant information is distributed across long, dense, and irregular histories that quickly exceed practical context limits.
To address current limitations, we introduce EHR2Path, a pathway-centric approach for forecasting and simulating full in-hospital patient pathways from routine EHRs. EHR2Path restructures multimodal hospital data, including structured events, laboratory measurements, medications, chart data, and clinical notes, into a unified, structured textual representation that preserves temporal structure within and across hospital stages (ED, ward, ICU). EHR2Path makes three key advances: (1) it performs pathway-level state forecasting and simulation by generating the next-hour EHR state over a broad set of clinical events and physiological values, rather than targeting a narrow endpoint; (2) it unifies the full breadth of EHR inputs, including structured, continuous, and free-text data, across ED, ward, and ICU in a single trajectory model; and (3) it introduces a Masked Summarization Bottleneck that restricts access to extended history through a small set of learned, section-specific summary tokens, while preserving recent events as uncompressed text, enabling efficient use of long clinical context. Given a patient’s past trajectory, EHR2Path predicts the next hour’s sparse EHR state and can be rolled out iteratively to simulate future pathways.
In a retrospective evaluation on the MIMIC-IV dataset (Johnson et al., 2023b), we show that EHR2Path improves next-hour forecasting and multi-day simulation, outperforming baselines in both forecasting and outcome prediction tasks while using context more efficiently. Together, these results establish a scalable path toward pathway-level modeling of heterogeneous in-hospital data for improved patient-state forecasting, simulation, and clinical decision support.
@@ 2 Related Work
The integration of heterogeneous EHR data into machine learning models has been a significant research focus, particularly with the adoption of sequence models like transformers (Vaswani, 2017). Existing methods fall into two categories: Outcome Prediction Models, that target specific patient outcomes, such as mortality, length-of-stay or diagnosis prediction, and Patient Timeline Prediction Models, which forecast full health trajectories but remain under-explored.
@@
Outcome Prediction Methods   Most EHR models predict outcomes using structured data like medical codes. Models such as BEHRT (Li et al., 2020), Med-BERT (Rasmy et al., 2021), CEHR-BERT (Pang et al., 2021), TransformEHR (Yang et al., 2023) have enhanced accuracy with temporal embeddings, artificial time tokens and multi-task learning. EHRSHOT (Wornow et al., 2023) introduced a benchmark integrating structured demographics and various coded events. Some other works include additional numerical or categorical data (Li et al., 2022; Pellegrini et al., 2023; Lovón-Melgarejo et al., 2024). Another major challenge in EHR modeling is Long-Context Integration, due to EHR’s time span and volume. Some outcome prediction models like Hi-BEHRT (Li et al., 2022), EhrMamba (Fallahpour et al., 2024), and CONTEXT (Wornow et al., 2024) extend token limits with hierarchical or sub-quadratic architectures, while models like REMed (Kim et al., 2024) and EMERGE (Zhu et al., 2024) propose retrieval-augmented approaches for selecting relevant features. UniHPF and GenHPF (Hur et al., 2022; 2023) show the effectiveness of combining structured and unstructured data into a unified text-based event representation with promising results on outcome classification tasks. Despite these advances, outcome prediction models typically target rather narrow downstream tasks and operate on a limited feature space, often focusing only on a fixed set of structured medical codes or few continuous signals, excluding a large set of available EHR data. They generally do not model or generate full patient trajectories.
@@
Patient Timeline Prediction Methods   In contrast to the prevalent outcome prediction methods, patient timeline prediction seeks to model and predict the entire sequence of a patient’s health events, offering a more holistic view of health trajectories. For instance, ESGPT (McDermott et al., 2023) provides a flexible library for transformer-based modeling of continuous-time event streams, including pre-processing utilities and example architectures. Recent approaches like CEHR-GPT (Pang et al., 2024) adapt GPT models for synthetic EHR generation, using structured tokens enriched with lab results, temporal and demographic information, but do not target patient-specific clinical forecasting. MOTOR (Steinberg et al., 2023) trains on timestamped sequences of medical events for time-to-event prediction, targeting individual outcomes like death or lab tests. Foresight (Kraljevic et al., 2024) predicts biomedical concepts from clinical notes, converting text into structured sequences using the SNOMED ontology (El-Sappagh et al., 2018). ETHOS (Renc et al., 2024) models structured EHR data from MIMIC-IV, generating tokenized Patient Health Timelines (PHTs). Yet, it omits unstructured notes and ICU chart events, key sources of patient information, and is limited to a 2,048-token input window. While these methods represent important progress toward trajectory modeling, they often operate on constrained input lengths, exclude unstructured modalities such as free-text notes and dense time-series signals such as ICU chart events and do not assess long-term simulation capabilities.
@@
Our work shifts the focus to forecasting and evaluating on entire hospital pathways, moving beyond isolated outcome prediction towards a comprehensive understanding of patient development. Unlike prior methods, we comprehensively integrate all available structured and unstructured EHR data and leverage LLMs to understand this heterogeneous and noisy information in its native form. To address long-context challenges, we introduce a novel summary mechanism that enables efficient modeling over extended time horizons and complex patient histories.
@@ 3 Scalable Modeling of Health Trajectories
We propose a novel approach to predict longitudinal patient health trajectories by integrating heterogeneous EHR data into a structured, language-based representation, as shown in Figure  2. Using an LLM backbone, we leverage its contextual understanding to model complex health data. To capture extended time horizons, we introduce a Masked Summarization Bottleneck that efficiently compresses longitudinal data. Additionally, we introduce two fine-tuning strategies for specializing EHR2Path to specific tasks. Our method enables comprehensive trajectory simulation, supporting holistic patient modeling.
@@ 3.1 Data Representation
EHR data is inherently heterogeneous, including static and temporal data, numerical and categorical features, and unstructured text from various clinical settings. We develop a unified textual representation of this data preserving its original form, avoiding heavy pre-processing to embrace real-world noise and incompleteness. This enhances robustness and supports diverse inputs without feature-specific adjustments, thereby improving scalability and real-world applicability. Our method supports diverse data types and uses natural language for feature names and values, instead of medical codes, enabling semantic interpretation by an LLM. An exemplary data sample is shown in section˜A.3.
@@
Clinical Units: At the highest level, $D_{p,t}$ is partitioned into clinical units $U$ = $\{\text{ED},\text{Hospital},\text{ICU}\}$, representing the Emergency Department, Hospital, and Intensive Care Unit.
@@
State Attributes and Diagnosis Codes   We incorporate state attributes to mark significant changes in the patient’s trajectory. These include events such as discharge from the ED, admission to and discharge from the ICU and Hospital, as well as death. At the final time step $T_{p}$ of a stay, ICD categories (Slee, 1978) corresponding to the patient’s diagnosis are added.
Input and Output Preparation   Model input and output are constructed by generating textual representations following the hierarchical structure defined above. The input includes all available data for a time window between twt-w up to the current time-point $t$, where $w$ denotes the input window size. To integrate longer time windows, we introduce a Masked Summarization Bottleneck, which condenses the entire patient history into one or multiple embedded representations for each category consisting of $N$ summary tokens each, allowing a window size of $t$. The output is formatted similarly but exclusively contains data recorded at the subsequent time-point $t+1$. To enhance efficiency, the output is sparse, including only the features that were actively recorded at $t+1$.
@@ 3.2 Patient Trajectory Prediction
Building upon the structured, language-based EHR representation outlined in section˜3.1, we now detail our approach for predicting future patient states and clinical pathways.
@@ 3.2.1 Model Architecture and Training
EHR2Path-Text (E2P-T): Processes a text representation of static patient data and events within the most recent $w$ hours (e.g., [tw,t][t-w,t]), capturing detailed recent observations in a structured format. Unless otherwise noted, E2P-T denotes the 24-hour text-only model, where $w=24$; alternative text windows are written explicitly, e.g., E2P-T-1h.
EHR2Path-Summ (E2P-S): Relies on summary embeddings of the entire patient history, generated via the Masked Summarization Bottleneck (section˜3.2.2), efficiently integrating long-term context.
EHR2Path-Summ+Text (E2P-S+T): Combines the text representation of recent events covering the 24-hour context used in E2P-T and the full history summary, balancing detailed short-term data with comprehensive longitudinal context.
Section˜A.4 provides detailed implementation and training details.
@@ 3.2.2 Masked Summarization Bottleneck
For E2P-S and E2P-S+T, we propose a Masked Summarization Bottleneck that produces compact, task-relevant representations for each clinical unit $u$ and data category $c$ to handle extensive patient histories. Formally, let $\mathbf{X}_{t}=\{x_{1},x_{2},\dots,x_{n}\}$ be the $n$ token long description of a section in the patient’s record at time $t$. We append $m$ summary tokens $\mathbf{S}=\{s_{1},\dots,s_{m}\}$ and $o$ output tokens $\mathbf{Y}=\{y_{1},\dots,y_{o}\}$ describing the values at $t+1$ for the current section, forming the sequence $\mathbf{X}^{\text{full}}_{t}=\{\mathbf{X}_{t},\mathbf{S},\mathbf{Y}\}$. A custom attention mask $\mathbf{M}$ constrains how tokens $i$ can attend to tokens $j$:
@@
As visualized in  fig.˜3, this ensures that output tokens $\{y_{1},\dots,y_{o}\}$ can only attend to the summary tokens $\{s_{1},\dots,s_{m}\}$ and preceding output tokens, creating a bottleneck. Following the information bottleneck principle (Tishby et al., 1999), our approach derives a compact representation $\mathbf{S}$ that maximizes mutual information $I(\mathbf{S};\mathbf{z}_{t+1})$ with the next state $\mathbf{z}_{t+1}$ while constraining $I(\mathbf{S};\mathbf{Z}_{1:t})$ with the patient history $\mathbf{Z}_{1:t}$ subject to the capacity $m$. Consequently, the Summarization Module is driven to compress the most relevant information to predict future states into the summary tokens, while discarding uninformative or redundant details. During inference, the summary tokens provide a lightweight embedding of the entire patient history for each data category $c$ in $D_{p,t}$. By handling each section separately, this allows us to in most cases include the entire patient history. At inference, the input tokens $\{x_{1},\dots,x_{n}\}$ for each category $c$ are combined with the summary tokens $\{s_{1},\dots,s_{m}\}$, and a single forward pass computes their hidden states $\mathbf{H}=\{\mathbf{h}_{1},\dots,\mathbf{h}_{n+m}\}$. The hidden states of the summary tokens, $\mathbf{H}^{\text{summary}}=\{\mathbf{h}_{n+1},\dots,\mathbf{h}_{n+m}\}$ are the final output of the Summary Model. This enables efficient summarization of all sections in $D_{p,t}$ without separate models, thus reducing computational overhead. Unlike traditional auto-encoder approaches that reconstruct every detail of the input, we optimize summary tokens for forecasting the patient’s next state. This task-focused approach avoids the inefficiencies associated with generic reconstructions and produces more effective representations for patient pathway prediction.
Figure 3: Masked Summarization Bottleneck. Input tokens encode past observations, while summary tokens (<<SUM>>) compress key information. A custom attention mask ensures outputs attend only to summaries, forcing the model to encode relevant patient data into a compact representation.